OpenStudy (anonymous):

a stone is dropped from the top of a cliff. It hits the ground below after 3.25s. How high is the cliff?

TheSmartOne (thesmartone):

You would need to use the formula for free falling objects which is Distance=.5xgxt^2

TheSmartOne (thesmartone):

x is times

TheSmartOne (thesmartone):

So t=3.25 t^2=10.5625s^2 g=9.8 m/(s^2) So distance=.5x9.8m x 10.5625 distance=51.75625m

TheSmartOne (thesmartone):

@marb